OpenStudy (anonymous):

Solve √(3)/2 = 25.5/f

OpenStudy (anonymous):

Simplify*

OpenStudy (anonymous):

f=25.5 *2/√(3) = 51/√(3)

OpenStudy (anonymous):

can you show me how you did that?

OpenStudy (anonymous):

=51√(3) /3

OpenStudy (anonymous):

17 square root of 3

OpenStudy (anonymous):

by cross multiplication

OpenStudy (anonymous):

you move f to one side and the equation to the other. You can't have a radical in the denominator so you multiply by root 3 and then simplify

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{2} = \frac{25.5}{f}\]\[f (\sqrt{3}) = 2(25.5)\]\[f = \frac{51}{\sqrt{3}} \approx 29.4448\]